For the following, use resistance values R1 = 40 2, R2 = 35 2, R3 = 30 12, R4 = 25 2 , R5 = 20 12, R6 = 15 2, R7 = 102, and RDA ER HER & Giov

Please use mesh-current analysis to solve for ix
Answer #1

Yix 25 20 - f-0401-22 + 10v From the bul in loop O => 4066,-4!1s (1,-14) +55 (9,- 12) = 0 90 °, - 35 12 -15 14=320 - 0 loop ③

By solving the above four equations i = 6.4.765. Amp iz = 5.4168 Amp iz = 4.3572 Amp Py = 4.8865 Amp ix- , -iz = 6.4765- 5.41

If you have any doubts please comment

If everything is clear please rate the answer

Thank you